# DAmodel: Hierarchical Bayesian Modelling of DA White Dwarfs for Spectrophotometric Calibration

### v1

@@ Abstract
We use hierarchical Bayesian modelling to calibrate a network of 32 all-sky faint DA white dwarf (DA WD) spectrophotometric standards ($19.516.5<V<$ 19.5) alongside the three CALSPEC standards, from 912 Å to 32 $\mu$m. The framework is the first of its kind to jointly infer photometric zeropoints and WD parameters ($\log g$, $T_{\text{eff}}$, $A_{V}$, $R_{V}$) by simultaneously modelling both photometric and spectroscopic data. We model panchromatic HST/WFC3 UVIS and IR fluxes, HST/STIS UV spectroscopy and ground-based optical spectroscopy to sub-percent precision. Photometric residuals for the sample are the lowest yet yielding $<0.004$ mag RMS on average from the UV to the NIR, achieved by jointly inferring time-dependent changes in system sensitivity and WFC3/IR count-rate nonlinearity. Our GPU-accelerated implementation enables efficient sampling via Hamiltonian Monte Carlo, critical for exploring the high-dimensional posterior space. The hierarchical nature of the model enables population analysis of intrinsic WD and dust parameters. Inferred SEDs from this model will be essential for calibrating the James Webb Space Telescope as well as next-generation surveys, including Vera Rubin Observatory’s Legacy Survey of Space and Time, and the Nancy Grace Roman Space Telescope.
@@ 1 Introduction
There is a particularly strong degeneracy between the photometric redshift of a galaxy and the relative calibration of a survey’s photometric zeropoints (High et al., 2009; Euclid Collaboration et al., 2024). If a survey’s relative zeropoints are calibrated incorrectly, the difference in flux could be misinterpreted as additional colour that would then compromise the photometric redshift estimate. This effect can also influence the Hubble diagram in SN Ia cosmology. Although systematic effects only change colours of individual objects by the order of a percent (Betoule et al., 2013), the combined effect of incorrectly measuring an entire SN population’s colour would have a significant influence on cosmological constraints (Hounsell et al., 2018). Scolnic et al. (2014) showed that colour calibration systematics can create an additional signal on the Hubble diagram at the 0.01-0.015 magnitude level. In the DESI analysis (DESI Collaboration et al., 2024), the difference between an evolving $w(z)$ and constant $w=-1$ dark energy equation of state Universe was a signal of only 0.04 mag, emphasising the importance of having confidence in relative photometric calibration. A further process in supernova cosmology that requires accurate relative flux calibration is the comparison of low redshift and high redshift samples. These contrasting samples are necessary to measure the change in the relative curve of the Hubble diagram, however, if these samples use inconsistent definitions of brightness or colour then they will be at odds with our true underlying cosmology (Efstathiou, 2024; Dhawan et al., 2024). To overcome this issue, it is imperative to perform reliable cross-calibration between surveys that rely on accurate flux standards.
@@
Spectrophotometric calibration relies on the precise modelling of the spectral energy distributions (SEDs) of astronomical objects that are well understood. Pure-hydrogen atmosphere DA white dwarfs (DA WDs) are ideal candidates to be modelled as standards, since they are near-perfect black bodies, with the addition of strong absorption features that are correlated with surface gravity and effective temperature (Holberg et al., 1985; Holberg & Bergeron, 2006; Hubeny & Lanz, 2017). So far, the three most accurate DA WDs used for spectrophotometric calibration have been GD71, GD153 and G191B2B. These WDs have little reddening and have been staples of the CALSPEC system (Bohlin et al., 2014, 2020), allowing for the calibration of astronomy’s most successful space missions, such as the Hubble Space Telescope (HST) and Gaia (Gaia Collaboration et al., 2016), as well as various other ground-based instruments. Since different missions scan different regions of the sky, it is necessary for networks of standards to have coverage over large portions of the celestial sphere. Rubin et al. (2022a) was awarded 35 HST/STIS orbits to observe 11 DA WDs to be added to CALSPEC system. Elms et al. (2024) recently expanded the CALSPEC standards by adding 17 new cool ($T_{\text{eff}}$ < 20,000 K) DA WDs. Low temperature DA WDs are particularly useful for standardisation as they have reduced non-local thermal equilibrium effects (more predictable SEDs) and are brighter in the near-infrared (NIR).
@@
An essential factor in spectrophotometric calibration that must be treated with care is the modelling of the effects of interstellar dust within the Milky Way. Dust along the line of sight between an object and the Earth has the combined effect of extinguishing and reddening light (Draine, 2003). Extinction, often measured in the $V$ band as $A_{V}$, is dependent on the column density of dust that lies between object and observer (e.g. Foster et al., 2013), which is also correlated with distance. The quantity $R_{V}=A_{V}/E(B-V)$, where $E(B-V)$ is the colour excess between $B$ band and $V$ band, is used to determine the amount of reddening one would expect from different types of dust grains. There have been various functional forms of dust relations that use these parameters to define extinction as a continuous function of wavelength $A(\lambda;A_{V},R_{V})$ (e.g. Cardelli et al., 1989; Fitzpatrick, 1999; Fitzpatrick et al., 2019; Gordon et al., 2023), that have been a key ingredient in spectrophotometric modelling DA WDs. Since DA WDs are so well understood, they are also useful tools to study dust. Sahu et al. (2024) recently used DA WDs to construct three-dimensional dust maps in the UV.
@@
The value of $A_{V}$ has a strong impact in the UV wavelengths, particularly around the 2175 Å bump that leads to a sharp drop off in extinction (Fitzpatrick & Massa, 1986). For this reason, it is important to spectroscopically observe DA WDs in the UV to constrain $A_{V}$ and extrapolate SEDs with a higher degree of confidence to longer wavelengths in the thermal and far infrared. Dust extinction decreases significantly in the NIR, meaning that if we have data coverage at these wavelengths we can disentangle degenerate dust and intrinsic parameters. The difference in extinction at UV and NIR wavelengths allows for more reliable inference of the $R_{V}$ parameters. This panchromatic strategy to mitigate dust effects is also used elsewhere in astronomy, for example in SN Ia modelling (Mandel et al., 2022; Thorp & Mandel, 2022; Grayling et al., 2024; Thorp et al., 2024a; Grayling & Popovic, 2024) and inference of galaxy properties Alsing et al. (2024); Thorp et al. (2024b).
@@
As we look towards the next-generation of deep surveys such as LSST (Ivezić et al., 2019), the Nancy Grace Roman Space Telescope (Roman; Green et al., 2012) and the James Webb Space Telescope (JWST; Gardner et al., 2006), we need dimmer standards to calibrate them. The three primary CALSPEC DA WDs, with magnitudes ranging from 11.7 < $V$ < 13.4, have been observed with JWST but are close to its observational limit and exceed the normal observing mode saturation limits of upcoming surveys, motivating the need for a new set of fainter standards. Our collaboration (led by Co-PIs Abhijit Saha and Gautham Narayan) has established an all-sky network of 32 faint (16.5 < $V$ < 19.5) DA white dwarfs to be used as spectrophotometric standards (Narayan et al., 2016, 2019; Calamida et al., 2019, 2022b; Axelrod et al., 2023). This group of DA WDs, illustrated in Figure 1, is intended to deliver exceptionally precise flux calibration, ready for the upcoming era of deep-sky surveys.
@@
The work of Narayan et al. (2019) was particularly successful in the optical and UV wavelengths, yielding an average RMS of 4 mmag across the five bands. The NIR F160W band proved to be more difficult due to known count-rate nonlinearity (CRNL) systematics (Riess et al., 2019; Bohlin & Deustua, 2019). Axelrod et al. (2023) subsequently expanded and reanalysed the network by using the approach developed by Narayan et al. (2019) to establish spectrophotometric priors on the model parameters. They then transitioned to a hierarchical Bayesian model to address the CRNL more accurately than using photometry alone. This combination proved successful, accomplishing a RMS of 8 mmag in the NIR, demonstrating the benefit of hierarchically analysing the network. The hierarchical approach enabled the inference of a band-dependent zeropoint shift relative to the CALSPEC system, which can only be derived from a set of stars. This allowed relative independence from CALSPEC which was crucial for ensuring the network could be calibrated to any magnitude system. It became clear that the next objective for this network would be a fully hierarchical end-to-end Bayesian analysis that would combine all spectroscopic and photometric data available.
@@ 2 Data
In Section 2.1 we describe how the standards were selected. In Section 2.2 and Section 2.3 we briefly describe the data previously analysed in Narayan et al. (2019) and Axelrod et al. (2023). More in-depth detail on the dataset can be found in Calamida et al. (2019) and Calamida et al. (2022b). Finally, in 2.4 we outline the data collection and reduction used for the new HST/STIS UV spectra.
@@ 2.1 Target Selection
Probable DA WDs in the northern and equatorial regions were selected by Narayan et al. (2019) and Calamida et al. (2019) using known sources, including Sloan Digital Sky Survey (SDSS) observations (Kleinman et al., 2004; Eisenstein et al., 2006) and the earlier Villanova catalogue (McCook & Sion, 1999), which has since been updated to the Montreal database (Dufour et al., 2017). These DA WD candidates had previously published low-resolution spectra, which provided enough detail to confirm them as DA type, assess their radiative properties, and check for any confounding factors such as magnetic line splitting or trace atmospheric elements. In contrast, no comprehensive list existed for faint DA WDs in the southern hemisphere, so candidates were identified from the SuperCOSMOS and VST surveys (Raddi et al., 2016, 2017) using photometric data and proper motion selection, focusing on stars with absolute magnitudes brighter than $V=$9.0. Only DA WDs with effective temperatures above 19,500 K were kept to be within the bounds where we have confidence in theoretical templates. It was also important to cut sources that are strongly influenced by magnetic fields, as these effects are not included in our modelling. This influence will present itself via Zeeman splitting of the Balmer lines, or in the presence of the Minkowski bands, shallow and broad absorption bands near 3650 Å, 4135 Å, and 4466 Å. Using the optical spectra, DA WDs were cut that exhibited strong magnetic traits, however, those with weaker effects could not be thrown away due to degeneracy with the surface gravity parameter.
The main goals of the study described in Calamida et al. (2022b) were to display spectra for all stars observed with the Southern Astrophysical Research (SOAR) telescope and to examine these stars for variability using the Las Cumbres Observatory (LCO) telescope network. Hot DA white dwarfs are not expected to vary intrinsically, but variability could arise due to binary companions, “seeing” effects from nearby faint red stars, or dust clouds surrounding the white dwarf. Narayan et al. (2019) and Calamida et al. (2019) eliminated four candidate stars based on spectroscopic and photometric data, while Calamida et al. (2022b) excluded a total of six stars from the full-sky set, leaving 32 faint stars and three brighter CALSPEC standards to establish our network. Figure 1 illustrates the locations and brightness of our faint all-sky network.
@@ 2.2 Hubble Space Telescope WFC3 Photometry
The optical photometry used in this study is extensively documented in Calamida et al. (2019) and Calamida et al. (2022b), while here we summarize the primary details of its collection. The photometry was acquired using the WFC3 with the UVIS and IR detectors, as part of Cycle 20 (average MJD=56452.94 days) and Cycle 22 (average MJD=57076.11 days), led by PI Abhijit Saha (proposals GO-12967 and GO-13711). Cycle 20 observations utilized five filters: F336W, F475W, F625W, F775W, and F160W. In Cycle 22, an additional near-UV (NUV) filter, F275W, was included to enhance detection of line-of-sight extinction. Some DA WD samples were only observed in one cycle, while others were observed in both. The latter group included nine DA WDs near the celestial equator. Observing this sample in both cycles not only improved photometric precision but also allowed for comparison between the two observing epochs.
Cycle 20 observations of the equatorial DA WDs were also used to validate the methodology of establishing spectrophotometric standards based on earlier-collected DA WDs (Narayan et al., 2016). Additionally, 14 candidate DA WDs and three CALSPEC WDs were observed during Cycle 22 (Bohlin et al., 2014, 2020), tying all observations to the HST photometric system. Collection methods in Cycle 20 varied by filter: F336W and F475W had three dithered exposures, while the other three filters had two exposures each. In Cycle 22, equatorial DA WDs were observed with three F275W exposures (with cosmic ray splits) and one additional exposure in each Cycle 20 filter. Ensuring at least three exposures per band for each source allowed for cross-exposure comparison, addressing issues like cosmic rays and detector artifacts in HST WFC3 images and maintaining sub-percent photometric precision. For WDs observed only in Cycle 22, five to seven dithered exposures per filter were obtained. The CALSPEC WDs were observed on three separate visits, with six to eight exposures per filter per visit. Observations in Cycle 20 used the UVIS1 sub-arrays and remained on the UVIS1 photometric system, while Cycle 22 observations used UVIS2 sub-arrays and were on UVIS2 photometric system.
The southern DA white dwarfs were observed with three HST Cycle 25 (average MJD=57588.79 days) orbits per star using WFC3 (Calamida et al., 2022b). The three main CALSPEC spectrophotometric standards GD71, GD153, and G191B2B each had three visits in Cycle 25 to address small potential sensitivity variations of WFC3. In 1.5 years, we expect only minor sensitivity variations, as sensitivity loss rates range from 0.1% to 0.2% per year, with greater losses at bluer wavelengths (Calamida et al., 2022a). Final photometric reductions were performed using Saha’s ILAPH code, as detailed in Calamida et al. (2019). This custom aperture photometry tool includes an interactive "growth curve" analysis for optimal sky subtraction. Ensuring the best possible count rates from the images was crucial, but equally important was ensuring that all measurements were placed on a consistent system of instrumental magnitudes. During post-processing, Cycle 25 observations were converted into the UVIS1 photometric system in the calwf3 calibration pipeline. Given that the HST/WFC3 data were obtained at different epochs over several years and with varying instrument configurations, special attention was given to adjusting for potential systematic magnitude shifts. The final photometry was placed on the AB magnitude system of CALSPEC. Although the use of AB magnitudes is not critical for subsequent analysis, it provided a convenient and self-consistent reference system. Ultimately, this process led to the determination of new AB magnitudes that are independent of the initial measurement system used for the observations.

### v2

@@ Abstract
We use hierarchical Bayesian modelling to calibrate a network of 32 all-sky faint DA white dwarf (DA WD) spectrophotometric standards ($19.516.5<V<$ 19.5) alongside three CALSPEC standards, from 912 Å to 32 $\mu$m. The framework is the first of its kind to jointly infer photometric zeropoints and WD parameters (surface gravity $\log g$, effective temperature $T_{\text{eff}}$, extinction $A_{V}$, dust relation parameter $R_{V}$) by simultaneously modelling both photometric and spectroscopic data. We model panchromatic Hubble Space Telescope Wide Field Camera 3 (HST/WFC3) UVIS and IR photometry, HST/STIS UV spectroscopy and ground-based optical spectroscopy to sub-percent precision. Photometric residuals for the sample are the lowest yet yielding $<0.004$ mag RMS on average from the UV to the NIR, achieved by jointly inferring time-dependent changes in system sensitivity and WFC3/IR count-rate nonlinearity. Our GPU-accelerated implementation enables efficient sampling via Hamiltonian Monte Carlo, critical for exploring the high-dimensional posterior space. The hierarchical nature of the model enables population analysis of intrinsic WD and dust parameters. Inferred spectral energy distributions from this model will be essential for calibrating the James Webb Space Telescope as well as next-generation surveys, including Vera Rubin Observatory’s Legacy Survey of Space and Time and the Nancy Grace Roman Space Telescope.
@@ 1 Introduction
There is a particularly strong degeneracy between the photometric redshift of a galaxy and the relative calibration of a survey’s photometric zeropoints (High et al., 2009; Euclid Collaboration et al., 2024). If a survey’s relative zeropoints are calibrated incorrectly, the difference in flux could be misinterpreted as additional colour that would then compromise the photometric redshift estimate. This effect can also influence the Hubble diagram in SN Ia cosmology. Although systematic effects only change colours of individual objects by the order of a percent (Betoule et al., 2013), the combined effect of incorrectly measuring an entire SN population’s colour would have a significant influence on cosmological constraints (Hounsell et al., 2018). Scolnic et al. (2014) showed that colour calibration systematics can create an additional signal on the Hubble diagram at the 0.01-0.015 magnitude level. In the DESI analysis (Adame et al., 2025), the difference between an evolving $w(z)$ and constant $w=-1$ dark energy equation of state universe was a signal of only 0.04 mag, emphasising the importance of having confidence in relative photometric calibration. A further process in supernova cosmology that requires accurate relative flux calibration is the comparison of low redshift and high redshift samples. These contrasting samples are necessary to measure the change in the relative curve of the Hubble diagram, however, if these samples use inconsistent definitions of brightness or colour then they will be at odds with our true underlying cosmology (Efstathiou, 2025; Dhawan et al., 2024). To overcome this issue, it is imperative to perform reliable cross-calibration between surveys that rely on accurate flux standards.
@@
Spectrophotometric calibration relies on the precise modelling of the spectral energy distributions (SEDs) of astronomical objects that are well understood. Pure-hydrogen atmosphere DA white dwarfs (DA WDs) are ideal candidates to be modelled as standards, since they have simple continuum-dominated spectra and have well characterised Balmer absorption features that are correlated with surface gravity and effective temperature (Holberg et al., 1985; Holberg & Bergeron, 2006; Hubeny & Lanz, 2017). So far, the three most accurate DA WDs used for spectrophotometric calibration have been GD71, GD153 and G191B2B. These WDs have little reddening and have been staples of the CALSPEC system (Bohlin et al., 2014, 2020), allowing for the calibration of astronomy’s most successful space missions, such as the Hubble Space Telescope (HST) and Gaia (Gaia Collaboration et al., 2016), as well as various other ground-based instruments. Since different missions scan different regions of the sky, it is necessary for networks of standards to have coverage over large portions of the celestial sphere. Rubin et al. (2022a) was awarded 35 HST/STIS orbits to observe 11 DA WDs to be added to CALSPEC system. Elms et al. (2024) recently expanded the CALSPEC standards by adding 17 new cool ($T_{\text{eff}}$ < 20 000 K) DA WDs. Low temperature DA WDs are particularly useful for standardisation as they have reduced non-local thermal equilibrium effects (more predictable SEDs) and are brighter in the near-infrared (NIR).
@@
An essential factor in spectrophotometric calibration that must be treated with care is the modelling of the effects of interstellar dust within the Milky Way. Dust along the line of sight between an object and the Earth has the combined effect of extinguishing and reddening light (Draine, 2003). Extinction, often measured in the $V$ band as $A_{V}$, is dependent on the column density of dust that lies between object and observer (e.g. Foster et al., 2013), which is also correlated with distance. The quantity $R_{V}=A_{V}/E(B-V)$, where $E(B-V)$ is the colour excess between $B$ band and $V$ band, is used to determine the amount of reddening one would expect from different types of dust grains. There have been various functional forms of dust relations that use these parameters to define extinction as a continuous function of wavelength $A(\lambda;A_{V},R_{V})$ (e.g. Cardelli et al., 1989; Fitzpatrick, 1999; Fitzpatrick et al., 2019; Gordon et al., 2021, 2023; Decleir et al., 2022), that have been a key ingredient in spectrophotometric modelling DA WDs. Since DA WDs are so well understood, they are also useful tools to study dust. Sahu et al. (2024) recently used DA WDs to validate three-dimensional dust maps in the UV.
@@
The value of $A_{V}$ has a strong impact in the UV wavelengths, particularly around the 2175 Å bump that leads to a sharp drop off in extinction (Fitzpatrick & Massa, 1986). For this reason, it is important to spectroscopically observe DA WDs in the UV to constrain $A_{V}$ and extrapolate SEDs with a higher degree of confidence to longer wavelengths in the thermal and far infrared. Dust extinction decreases significantly in the NIR, meaning that if we have data coverage at these wavelengths we can disentangle degenerate dust and intrinsic parameters. The difference in extinction at UV and NIR wavelengths allows for more reliable inference of the $R_{V}$ parameters. This panchromatic strategy to mitigate dust effects is also used elsewhere in astronomy, for example in modelling SNe Ia (Type Ia Supernovae) (Mandel et al., 2022; Thorp & Mandel, 2022; Grayling et al., 2024; Thorp et al., 2024a; Grayling & Popovic, 2024) and inference of galaxy properties Alsing et al. (2024); Thorp et al. (2024b).
@@
As we look towards the next-generation of deep surveys such as LSST (Ivezić et al., 2019), the Nancy Grace Roman Space Telescope (Roman; Green et al., 2012) and the James Webb Space Telescope (JWST; Gardner et al., 2006), we need dimmer standards to calibrate them. The three primary CALSPEC DA WDs, with magnitudes ranging from 11.7 < $V$ < 13.4, have been observed with JWST but are close to its observational limit and exceed the normal observing mode saturation limits of upcoming surveys, motivating the need for a new set of fainter standards. Our collaboration (initiated and led by co-author Saha) has established an all-sky network of 32 faint (16.5 < $V$ < 19.5) DA white dwarfs to be used as spectrophotometric standards (Narayan et al., 2016, 2019; Calamida et al., 2019, 2022b; Axelrod et al., 2023). This group of DA WDs, illustrated in Figure 1, is intended to deliver exceptionally precise flux calibration, ready for the upcoming era of deep-sky surveys.
@@
The work of Narayan et al. (2019) was particularly successful in the optical and UV wavelengths, yielding an average RMS of 4 mmag across the five bands. The NIR F160W band (centred at 1537 nm) proved to be more difficult due to known count-rate nonlinearity (CRNL) systematics (Riess et al., 2019; Bohlin & Deustua, 2019). Axelrod et al. (2023) subsequently expanded and reanalysed the network by using the approach developed by Narayan et al. (2019) to establish spectrophotometric priors on the model parameters. They then transitioned to a hierarchical Bayesian model to address the CRNL more accurately than using photometry alone. This combination proved successful, accomplishing a RMS of 8 mmag in the NIR, demonstrating the benefit of hierarchically analysing the network. The hierarchical approach enabled the inference of a band-dependent zeropoint shift relative to the CALSPEC system, which can only be derived from a set of stars. This allowed relative independence from CALSPEC which was crucial for ensuring the network could be calibrated to any magnitude system. It became clear that the next objective for this network would be a fully hierarchical end-to-end Bayesian analysis that would combine all spectroscopic and photometric data available.
@@ 2 Data
All the data used in this work has been previously described and analysed individually in other publications, however, this is the first time in which all the data has been modelled simultaneously. The northern and equatorial standards were analysed in Narayan et al. (2019) using optical spectra and photometry detailed in Calamida et al. (2019). The photometry and optical spectra of the full network was analysed in Axelrod et al. (2023) and was introduced in Calamida et al. (2022b). More recently, the HST UV spectra for a subset of the network was published and analysed in Bohlin et al. (2025). In this section we briefly describe the DA WD network and the observational data used in our work.
@@ 2.1 Target Selection
Narayan et al. (2019) and Calamida et al. (2019) selected 19 northern and equatorial faint DA WDs using known sources, while Calamida et al. (2022b) established an additional 13 southern standards using their own methodology. Standards were carefully selected to avoid for any factors such as magnetic line splitting or trace atmospheric elements in their spectra, as these effects were not included in our modelling. Standards were also monitored to ensure there was no photometric variability. Hot DA white dwarfs are not expected to vary intrinsically, but variability could arise due to binary companions, “seeing” effects from nearby faint red stars or dust clouds surrounding the white dwarf. Absolute brightnesses of the stars were chosen to be no brighter than $V=9.0$. Only DA WDs with effective temperatures above 19 500 K were kept to be within the bounds where we have confidence in theoretical templates. The final all-sky set of 32 faint DA WDs that met the selection criteria was published in Calamida et al. (2022b). These are combined with the three CALSPEC primary standards to make the full network of 35 stars that we analyse in our work. Effective temperatures of the network range from 20 000 K to 67 000 K and the natural logarithm of surface gravities are between 7.2 and 9.0. The Gaia $G$ band apparent brightnesses of the faint standards range between 15.9 mag and 20.0 mag. Figure 1 illustrates the locations and brightnesses of our faint all-sky network. Coordinates of the WDs can be found in Table 1 and online at https://zenodo.org/records/14939532/files/wdfs_coords.txt.
@@ 2.2 Hubble Space Telescope WFC3 Photometry
The optical photometry used in this study is extensively documented in Calamida et al. (2019) and Calamida et al. (2022b), while here we summarise important details for our modelling. The photometry was acquired with the WFC3 UVIS and IR channels during Cycle 20 (average MJD=56452.94 days), Cycle 22 (average MJD=57076.11 days) and Cycle 25 (average MJD=57588.79 days) with programs GO-12967, GO-13711 and GO-15113 respectively (PI Abhijit Saha). Cycle 20 observed six of the northern and equatorial faint standards in five HST filters: F336W, F475W, F625W, F775W, and F160W. Cycle 22 observed all 19 of the northern and equatorial faint standards, as well as the three CALSPEC primary standards, in the same five filters and the additional near-UV F275W filter. Cycle 25 observed the 13 southern standards and the three primary CALSPEC standards using all six bands from the near-UV to the near-infrared. The sensitivity of WFC3 has changed as a function of time (Calamida et al., 2022a), so we can use the differences of the faint standards observed in both Cycle 20 and 22, as well as change in the CALSPEC primary standards in Cycle 22 and 25, to help constrain the varying zeropoints.
The bands in the near-UV and optical were on different UVIS sub-arrays in each cycle. Observations in Cycle 20 used the UVIS1 sub-arrays and remained on the UVIS1 photometric system, while Cycle 22 observations used UVIS2 sub-arrays and were on UVIS2 photometric system. Observations in Cycle 25 were taken with the UVIS2 sub-arrays but converted back to the UVIS1 photometric system in the calwf3 calibration pipeline. The differences in photometric system are important to consider when modelling the changes in zeropoints between cycles. The final photometry was placed on the AB magnitude system of CALSPEC. Although the use of AB magnitudes is not critical for subsequent analysis, it provides a convenient and self-consistent reference system.